Here's my question.
There was a 10K race. 112 more men than women were entered in the race. The sum of men and women is 600. How many women were entered in the race?

Answer by oscargut(2103)

Let x number of men
y number of women
then x = y+112 (112 more men than women were entered in the race)
and x+y=600 (The sum of men and women is 600.)
then y+112+y=600 then 2y+112=600 then 2y=488 then y=244 and x=244+112=356
R: x=356 and y=244